OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

–2 1 –1 2 these are the answers givin

OpenStudy (anonymous):

m = (y2 - y1) / (x2 - x1) = (1 - - 4)/ (-2 - 3 ) = 5/-5 = -1

OpenStudy (anonymous):

thank you